My syllabus gives these co-ordinance as the solution but I don't understand how they got this answer?
2x + y = 17
y = 6x + 1
x = 2, y = 13
Thanks

Answer by mananth(16949)

2x + y = 17
y = 6x + 1

y= 2x+1
substitute y in the first equation
2x+(6x+1)=17
8x+1=17
-1 on both sides
8x+1-1=17-1
8x=16
/2
x=2
Plug x=2 in y=2x+1
y=6*2+1
y=13